OpenStudy (anonymous):

write the given equation in standard form: x^2+y^2-4x-6y-13=0

OpenStudy (anonymous):

How about (x-2)^2 + (y-3)^2 = 26 http://www.oojih.com for free study guide and review questions

OpenStudy (anonymous):

do you know how to 'complete the square"?

OpenStudy (anonymous):

first rewrite as \[x^2-4x+y^2-6x=13\] then half of -4 is -2 and half of -6 is -3 so write as \[(x-2)^2+(y-3)^2=13+2^2+3^2=13+4+9=26

OpenStudy (anonymous):

\[(x-2)^2+(y-3)^2=13+2^2+3^2=13+4+9=26\]

OpenStudy (watchmath):

\((x-2)^2+(y-3)^2=13+4+9\) \((x-2)^2+(y-2)^2=30\) \(\frac{(x-2)^2}{30}+\frac{(y-3)^2}{30}=1\)

OpenStudy (watchmath):

lol you were correct satellite

OpenStudy (anonymous):

hi watchman

OpenStudy (anonymous):

yeah arithmetic is tough this time of night!

OpenStudy (watchmath):

still energetic answering questions :)?

OpenStudy (anonymous):

so is typing "watchmath" instead of "watchman" sorry

OpenStudy (anonymous):

thanks guys, very helpful! I appreciate it!

OpenStudy (anonymous):

bb911 is it clear how to complete a square?

OpenStudy (anonymous):

yes, I understand now, thanks!